OpenStudy (anonymous):

if theres 76 questions and i get 60 right ,what percentage is that?

OpenStudy (anonymous):

78.9%

jimthompson5910 (jim_thompson5910):

Percentage = (number correct/number total)*100 Percentage = (60/76)*100 Percentage = (0.78947)*100 Percentage = 78.947%

OpenStudy (anonymous):

thats guys! :)